OpenStudy (anonymous):

A cubic polynomial function of degree f had leading coefficient 2 and constant term 7. If f(1)= 7 and f(2) = 9, what is f(-2)?

OpenStudy (anonymous):

Please help I have a test tomorrow!

jimthompson5910 (jim_thompson5910):

f(x) = ax^3 + bx^2 + cx + d f(x) = 2x^3 + bx^2 + cx + 7 f(1) = 2(1)^3 + b(1)^2 + c(1) + 7 f(1) = 2(1) + b(1) + c(1) + 7 f(1) = 2 + b + c + 7 f(1) = b + c + 9 7 = b + c + 9 7-9 = b + c -2 = b + c b+c = -2 ------------------------ f(x) = 2x^3 + bx^2 + cx + 7 f(2) = 2(2)^3 + b(2)^2 + c(2) + 7 f(2) = 2(8) + b(4) + c(2) + 7 f(2) = 16 + 4b + 2c + 7 f(2) = 4b + 2c + 23 9 = 4b + 2c + 23 9-23 = 4b + 2c -14 = 4b + 2c 4b + 2c = -14

jimthompson5910 (jim_thompson5910):

You now have the two equations b+c = -2 4b + 2c = -14 which you can use to solve for b and c Once you have the values of b and c, you can use them to find f(-2)

OpenStudy (anonymous):

oh my lord thank you so much seriously thank you

jimthompson5910 (jim_thompson5910):

np